# OneOcc: Semantic Occupancy Prediction for Legged Robots with a Single Panoramic Camera

### v1

@@ Abstract
Robust 3D semantic occupancy is crucial for legged/humanoid robots, yet most semantic scene completion (SSC) systems target wheeled platforms with forward-facing sensors. We present OneOcc, a vision-only panoramic SSC framework designed for gait-introduced body jitter and $360^{\circ}$ continuity. OneOcc combines: (i) Dual-Projection fusion (DP-ER) to exploit the annular panorama and its equirectangular unfolding, preserving $360^{\circ}$ continuity and grid alignment; (ii) Bi-Grid Voxelization (BGV) to reason in Cartesian and cylindrical-polar spaces, reducing discretization bias and sharpening free/occupied boundaries; (iii) a lightweight decoder with Hierarchical AMoE-3D for dynamic multi-scale fusion and better long-range/occlusion reasoning; and (iv) plug-and-play Gait Displacement Compensation (GDC) learning feature-level motion correction without extra sensors. We also release two panoramic occupancy benchmarks: QuadOcc (real quadruped, first-person $360^{\circ}$) and Human360Occ (H3O) (CARLA human-ego $360^{\circ}$ with RGB, Depth, semantic occupancy; standardized within-/cross-city splits). OneOcc sets new state-of-the-art (SOTA): on QuadOcc it beats strong vision baselines and popular LiDAR ones; on H3O it gains +3.83 mIoU (within-city) and +8.08 (cross-city). Modules are lightweight, enabling deployable full-surround perception for legged/humanoid robots. Datasets and code will be publicly available at https://github.com/MasterHow/OneOcc.
@@ 1 Introduction
Figure 1: Teaser. (a) Single-sensor panoramic semantic scene completion (SSC) on a quadruped: an omnidirectional camera observes full $360^{\circ}$ with gait-induced motion; OneOcc outputs voxelized semantic occupancy. (b) On QuadOcc, OneOcc reaches 20.56 mIoU, surpassing the best LiDAR baseline (LMSCNet, $18.44$) and vision baseline (MonoScene, $19.19$). On H3O, OneOcc attains 37.29/32.23 mIoU under within-/cross-city, outperforming the best vision baselines ($33.46/24.15$).
A holistic 3D understanding of nearby surroundings is fundamental for robotic autonomy in navigation [wang2024omega, hoeller2022neural, wang2024he_nav, wang2024agrnav, patel2025hierarchical, li2024stereonavnet], locomotion [niijima2025realtime_multi_plane, dong2025marg, escontrela2025gaussgym, chen2025learning_traversal], manipulation [kartmann2021semantic, dengler2025efficient_manipulation, li2025manipdreamer3d, wang2025odyssey], and human-robot interaction [wu2025embodiedocc, wang2025embodiedocc++, liu2024revisit_human_scene_interaction]. Among 3D perception tasks, Semantic Scene Completion (SSC)—predicting complete volumetric geometry and semantics from partial observations—has become central [lai2024rtonet, cao2024slcf_sequential, yan2024pointssc, zhang2025roboocc], with most progress driven by autonomous driving where forward-facing sensors on stable wheeled platforms yield benign motion and sensing.
@@
NYUv2 [nyu] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 1.4K 1.4K [Uncaptioned image]
ScanNet [scannet] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(31{\times}62{\times}62)$ 1.5K 1.5K [Uncaptioned image]
SUNCG [song2017semantic] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 46K 140K [Uncaptioned image]
SemanticPOSS [semanticposs] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] L $(32{\times}256{\times}256)$ 6 3K [Uncaptioned image]
SemanticKITTI [semantickitti] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 22 9K [Uncaptioned image]
nuScenes [wang2023openoccupancy] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(40{\times}512{\times}512)$ 850 200K [Uncaptioned image]
Occ3D-Waymo [tian2024occ3d] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(32{\times}200{\times}200)$ 1K 200K [Uncaptioned image]
SSCBench-KITTI-360 [li2024sscbench] Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}200{\times}200)$ 11 91K [Uncaptioned image]
ScribbleSC [wang2024label_efficient] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 11 9K [Uncaptioned image]
OmniHD-Scenes [zheng2024omnihd] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&R 200 60K [Uncaptioned image][Uncaptioned image]
WildOcc [zhai2024wildocc] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(40{\times}100{\times}100)$ 5 10K [Uncaptioned image]
DSEC-SSC [guo2025event] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&E $(16{\times}128{\times}128)$ 6 3K [Uncaptioned image]
ORAD-3D [min2025advancing_orad] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L 145 58K [Uncaptioned image]
Co3SOP [wu2025synthetic_v2x] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(70{\times}1000{\times}1000)$ [Uncaptioned image]
@@
Panoramic SSC with a single sensor. A single panoramic camera provides compact, true $360^{\circ}$ coverage but introduces annular distortions and seam artifacts after unwrapping [lin2025one_flight]—issues poorly handled by perspective-camera SSC pipelines. This motivates a design that preserves raw panoramic cues while exploiting grid-friendly projections.
@@
Benchmarks: QuadOcc and Human360Occ (H3O). To advance panoramic SSC on legged/humanoid robots, we release two complementary benchmarks. QuadOcc is a real first-person $360^{\circ}$ dataset on a quadruped within a campus domain, with standardized day/dusk/night coverage and semi-automatic ground truth (multi-frame LiDAR aggregation, Grounded-SAM [ren2024grounded] initialization, targeted manual fixes). It contains 10 scenes and 24K frames, uses 6 semantic categories for training. The occupancy uses 0.4 m voxels on a $64\!\times\!64\!\times\!8$ grid—reflecting embodied compute and lower speed than cars. Human360Occ (H3O) is CARLA-based human-ego $360^{\circ}$ with simulated gait, 160 sequences / 8K frames over 16 maps and multiple weathers/times; each frame provides RGB, metric depth, semantic occupancy at $\times64$$\times$8 and $\times128$$\times$16, and pose. It supports within-city (per-map 8:28{:}2 train/val) and cross-city (default 12/412{/}4 maps) splits.
@@ 2 Related Work
Semantic scene understanding for legged robots. Legged autonomy motivates perception tailored to agile motion and strict payload budgets. Prior efforts include vision-based terrain reconstruction [zhang2022vision], large-scale datasets [patel2025tartanground], and traversability learning [deng2021vision, aegidius2025watch, oh2024trip]; complementary modules cover robust tracking [cao2025siamese, xin2024robotic] and instruction grounding [mei2024quadrupedgpt, han2025space]. With increasing evidence that limited FoV hinders agility [li2024move, cheng2024quadruped, wang2025omni], panoramic sensors are adopted for omnidirectional tracking [luo2025omnidirectional], multimodal odometry [li2025limo], and even satellite communication on the move [liu2025starlink]. Some works study probabilistic semantic mapping [chen2025particle], and humanoid systems explore panoramic/occupancy perception [cui2025humanoid, zhang2025humanoidpano, zhang2025roboocc, zhang2025occupancy_world_models_robots]. However, dense 3D semantic occupancy from a single panoramic camera on dynamic legged platforms is still largely open. We address this by releasing QuadOcc and Human360Occ and proposing OneOcc, a lightweight panoramic occupancy framework built around dual-projection fusion, bi-grid voxelization, hierarchical 3D MoE fusion, and gait displacement compensation.
@@ 3.3 Unwrapping and Dual-Projection Encoders
We calibrate the omnidirectional panoramic annular lens (PAL) camera with the Taylor polynomial model [scaramuzza2006toolbox]. The unwrapping samples the raw annulus at coordinates given by the calibrated inverse mapping:
@@
SSCNet [song2017semantic] L 0.00 0.04 44.34 16.06 22.41 4.77 63.42 65.82 47.71 14.60
SSCNet-full [song2017semantic] L 0.71 0.04 55.14 14.93 19.66 7.76 78.81 60.03 51.69 16.38
LMSCNet [roldao2020lmscnet] L 0.88 0.20 57.02 16.45 24.39 11.70 79.70 64.49 55.40 18.44
OccRWKV [wang2024occrwkv] L 2.05 0.35 54.17 10.29 24.92 8.38 71.98 62.80 50.47 16.69
@@
VoxFormer-S [li2023voxformer] P+D 0.31 1.67 36.83 6.48 7.09 3.81 45.96 69.95 38.38 9.36
SGN [mei2023camera] P+D 7.39 1.88 53.35 9.06 21.95 9.39 59.23 67.41 46.05 17.17
VoxFormer-S [li2023voxformer] P+L 0.24 1.66 38.51 10.90 11.88 4.45 51.29 75.84 44.08 11.27
VoxFormer-T [li2023voxformer] P+L 0.30 1.65 40.21 14.48 15.86 4.64 56.94 77.10 48.70 12.86
SGN [mei2023camera] P+L 11.62 2.47 53.06 15.60 25.67 9.91 73.70 59.62 49.16 19.72
OccFormer [zhang2023occformer] P 0.29 0.37 49.46 10.36 15.00 2.64 45.79 76.99 40.28 13.02
MonoScene [cao2022monoscene] P 8.15 1.59 55.66 12.88 26.10 10.78 62.10 69.28 48.69 19.19
@@
VoxFormer-[li2023voxformer] P+Dpred. 22.27 17.44 2.86 5.65 0.87 0.14 0.00 0.00 61.46 0.18 40.21 59.49 31.57 11.09
VoxFormer-[li2023voxformer] P+Dgt 35.84 25.86 12.32 9.90 3.73 0.70 0.19 0.17 64.79 0.93 59.91 71.12 48.19 15.44
SGN-[mei2023camera] P+Dpred. 54.55 44.61 15.51 35.16 23.91 7.68 4.52 15.77 65.24 19.46 61.82 64.36 46.05 28.64
SGN-[mei2023camera] P+Dpred. 54.81 47.44 15.64 33.22 24.23 8.53 2.62 13.51 65.88 19.72 62.28 63.72 45.98 28.56
OccFormer [zhang2023occformer] P 54.11 46.41 15.14 32.84 17.57 4.30 0.00 1.65 65.08 11.72 57.30 69.94 45.98 24.88
MonoScene [cao2022monoscene] P 62.45 57.18 19.97 38.91 29.76 13.77 5.03 16.76 70.15 20.67 68.78 71.66 54.08 33.46
@@
VoxFormer-[li2023voxformer] P+Dpred. 19.02 15.82 2.84 5.86 0.95 0.10 0.00 0.01 61.62 0.10 36.68 52.00 27.40 10.63
VoxFormer-[li2023voxformer] P+Dgt 32.55 25.95 8.98 9.93 4.68 0.44 0.00 0.00 65.45 0.26 59.08 61.21 42.99 14.82
SGN-[mei2023camera] P+Dpred. 42.57 23.94 9.42 15.84 18.22 10.21 1.55 3.11 63.44 11.93 45.84 47.11 30.26 20.02
SGN-[mei2023camera] P+Dpred. 43.47 24.70 9.19 20.27 17.21 8.21 1.85 5.08 61.10 14.89 47.82 50.98 32.76 20.60
OccFormer [zhang2023occformer] P 47.00 39.01 10.14 24.85 11.84 1.52 0.05 0.77 66.60 6.93 53.12 63.57 40.73 20.87
MonoScene [cao2022monoscene] P 49.68 41.41 11.04 20.57 20.34 10.36 1.98 2.79 69.68 13.69 67.39 55.00 43.44 24.15

### v2

@@ Abstract
Robust 3D semantic occupancy is crucial for legged/humanoid robots, yet most semantic scene completion (SSC) systems target wheeled platforms with forward-facing sensors. We present OneOcc, a vision-only panoramic SSC framework designed for gait-introduced body jitter and $360^{\circ}$ continuity. OneOcc combines: (i) Dual-Projection fusion (DP-ER) to exploit the annular panorama and its equirectangular unfolding, preserving $360^{\circ}$ continuity and grid alignment; (ii) Bi-Grid Voxelization (BGV) to reason in Cartesian and cylindrical-polar spaces, reducing discretization bias and sharpening free/occupied boundaries; (iii) a lightweight decoder with Hierarchical AMoE-3D for dynamic multi-scale fusion and better long-range/occlusion reasoning; and (iv) plug-and-play Gait Displacement Compensation (GDC) learning feature-level motion correction without extra sensors. We also release two panoramic occupancy benchmarks: QuadOcc (real quadruped, first-person $360^{\circ}$) and Human360Occ (H3O) (CARLA human-ego $360^{\circ}$ with RGB, Depth, semantic occupancy; standardized within-/cross-city splits). OneOcc sets a new state of the art on QuadOcc, outperforming strong vision baselines and remaining competitive with classical LiDAR baselines; on H3O it gains +3.83 mIoU (within-city) and +8.08 (cross-city). Modules are lightweight, enabling deployable full-surround perception for legged/humanoid robots. Datasets and code will be publicly available at OneOcc.
@@ 1 Introduction
Figure 1: Teaser. (a) Single-sensor panoramic semantic scene completion (SSC) on a quadruped: an omnidirectional camera observes full $360^{\circ}$ with gait-induced motion; OneOcc outputs voxelized semantic occupancy. (b) Summary on QuadOcc and Human360Occ (H3O): bars show mIoU(%). On QuadOcc, OneOcc reaches 20.56 mIoU, surpassing the best LiDAR baseline (LMSCNet, $18.44$) and vision baseline (MonoScene, $19.19$). On H3O, OneOcc attains 37.29/32.23 mIoU under within-/cross-city, outperforming the best vision baselines ($33.46/24.15$).
A holistic 3D understanding of nearby surroundings is fundamental for robotic autonomy in navigation [82, 28, 83, 84, 67, 39], locomotion [61, 19, 21, 11], manipulation [33, 18, 44, 86], and human-robot interaction [93, 81, 48]. Among 3D perception tasks, Semantic Scene Completion (SSC)—predicting complete volumetric geometry and semantics from partial observations—has become central [36, 5, 99, 108], with most progress driven by autonomous driving where forward-facing sensors on stable wheeled platforms yield benign motion and sensing.
@@
NYUv2 [75] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 1.4K 1.4K [Uncaptioned image]
ScanNet [16] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(31{\times}62{\times}62)$ 1.5K 1.5K [Uncaptioned image]
SUNCG [76] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Indoor [Uncaptioned image] [Uncaptioned image] C&D $(144{\times}240{\times}240)$ 46K 140K [Uncaptioned image]
SemanticPOSS [66] n.a. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] L $(32{\times}256{\times}256)$ 6 3K [Uncaptioned image]
SemanticKITTI [2] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 22 9K [Uncaptioned image]
nuScenes [89] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(40{\times}512{\times}512)$ 850 200K [Uncaptioned image]
Occ3D-Waymo [80] Pinhole+Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L $(32{\times}200{\times}200)$ 1K 200K [Uncaptioned image]
SSCBench-KITTI-360 [42] Fisheye [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}200{\times}200)$ 11 91K [Uncaptioned image]
ScribbleSC [88] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(32{\times}256{\times}256)$ 11 9K [Uncaptioned image]
OmniHD-Scenes [112] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&R 200 60K [Uncaptioned image][Uncaptioned image]
WildOcc [103] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(40{\times}100{\times}100)$ 5 10K [Uncaptioned image]
DSEC-SSC [26] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image] C&L&E $(16{\times}128{\times}128)$ 6 3K [Uncaptioned image]
ORAD-3D [58] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image][Uncaptioned image][Uncaptioned image] C&L 145 58K [Uncaptioned image]
Co3SOP [92] Pinhole [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Outdoor [Uncaptioned image] [Uncaptioned image] C&L $(70{\times}1000{\times}1000)$ [Uncaptioned image]
@@
Panoramic SSC with a single sensor. A single panoramic camera provides compact, true $360^{\circ}$ coverage but introduces annular distortions and seam artifacts after unwrapping [45]—issues poorly handled by perspective-camera SSC pipelines. This motivates a design that preserves raw panoramic cues while exploiting grid-friendly projections.
@@
Benchmarks: QuadOcc and Human360Occ (H3O). To advance panoramic SSC on legged/humanoid robots, we release two complementary benchmarks. QuadOcc is a real first-person $360^{\circ}$ dataset on a quadruped within a campus domain, with standardized day/dusk/night coverage and semi-automatic ground truth (multi-frame LiDAR aggregation, Grounded-SAM [70] initialization, targeted manual fixes). It contains 10 scenes and 24K frames, uses 6 semantic categories for training. The occupancy uses 0.4 m voxels on a $64\!\times\!64\!\times\!8$ grid—reflecting embodied compute and lower speed than cars. Human360Occ (H3O) is CARLA-based human-ego $360^{\circ}$ with simulated gait, 160 sequences / 8K frames over 16 maps and multiple weathers/times; each frame provides RGB, metric depth, semantic occupancy at $\times64$$\times$8 and $\times128$$\times$16, and pose. It supports within-city (per-map 8:28{:}2 train/val) and cross-city (default 12/412{/}4 maps) splits. See details on the construction pipeline and dataset statistics in the supplementary material.
@@ 2 Related Work
Semantic scene understanding for legged robots. Legged autonomy motivates perception tailored to agile motion and strict payload budgets. Prior efforts include vision-based terrain reconstruction [104], large-scale datasets [68], and traversability learning [17, 1, 62]; complementary modules cover robust tracking [6, 95] and instruction grounding [57, 27]. With increasing evidence that limited FoV hinders agility [41, 13, 90], panoramic sensors are adopted for omnidirectional tracking [52], multimodal odometry [40], and even satellite communication on the move [46]. Some works study probabilistic semantic mapping [9], and humanoid systems explore panoramic/occupancy perception [15, 106, 108, 109]. However, dense 3D semantic occupancy from a single panoramic camera on dynamic legged platforms is still largely open. We address this by releasing QuadOcc and Human360Occ and proposing OneOcc, a lightweight panoramic occupancy framework built around dual-projection fusion, bi-grid voxelization, hierarchical 3D MoE fusion, and gait displacement compensation.
@@ 3.3 Unwrapping and Dual-Projection Encoders
We calibrate the omnidirectional panoramic annular lens (PAL) camera with the Taylor polynomial model [72]. The unwrapping samples the raw annulus at coordinates given by the calibrated inverse mapping:
@@
SSCNet [76] L 0.00 0.04 44.34 16.06 22.41 4.77 63.42 65.82 47.71 14.60
SSCNet-full [76] L 0.71 0.04 55.14 14.93 19.66 7.76 78.81 60.03 51.69 16.38
LMSCNet [71] L 0.88 0.20 57.02 16.45 24.39 11.70 79.70 64.49 55.40 18.44
OccRWKV [85] L 2.05 0.35 54.17 10.29 24.92 8.38 71.98 62.80 50.47 16.69
@@
VoxFormer-S [43] P+D 0.31 1.67 36.83 6.48 7.09 3.81 45.96 69.95 38.38 9.36
SGN [56] P+D 7.39 1.88 53.35 9.06 21.95 9.39 59.23 67.41 46.05 17.17
VoxFormer-S [43] P+L 0.24 1.66 38.51 10.90 11.88 4.45 51.29 75.84 44.08 11.27
VoxFormer-T [43] P+L 0.30 1.65 40.21 14.48 15.86 4.64 56.94 77.10 48.70 12.86
SGN [56] P+L 11.62 2.47 53.06 15.60 25.67 9.91 73.70 59.62 49.16 19.72
OccFormer [107] P 0.29 0.37 49.46 10.36 15.00 2.64 45.79 76.99 40.28 13.02
MonoScene [4] P 8.15 1.59 55.66 12.88 26.10 10.78 62.10 69.28 48.69 19.19
@@
VoxFormer-[43] P+Dpred. 22.27 17.44 2.86 5.65 0.87 0.14 0.00 0.00 61.46 0.18 40.21 59.49 31.57 11.09
VoxFormer-[43] P+Dgt 35.84 25.86 12.32 9.90 3.73 0.70 0.19 0.17 64.79 0.93 59.91 71.12 48.19 15.44
SGN-[56] P+Dpred. 54.55 44.61 15.51 35.16 23.91 7.68 4.52 15.77 65.24 19.46 61.82 64.36 46.05 28.64
SGN-[56] P+Dpred. 54.81 47.44 15.64 33.22 24.23 8.53 2.62 13.51 65.88 19.72 62.28 63.72 45.98 28.56
OccFormer [107] P 54.11 46.41 15.14 32.84 17.57 4.30 0.00 1.65 65.08 11.72 57.30 69.94 45.98 24.88
MonoScene [4] P 62.45 57.18 19.97 38.91 29.76 13.77 5.03 16.76 70.15 20.67 68.78 71.66 54.08 33.46
@@
VoxFormer-[43] P+Dpred. 19.02 15.82 2.84 5.86 0.95 0.10 0.00 0.01 61.62 0.10 36.68 52.00 27.40 10.63
VoxFormer-[43] P+Dgt 32.55 25.95 8.98 9.93 4.68 0.44 0.00 0.00 65.45 0.26 59.08 61.21 42.99 14.82
SGN-[56] P+Dpred. 42.57 23.94 9.42 15.84 18.22 10.21 1.55 3.11 63.44 11.93 45.84 47.11 30.26 20.02
SGN-[56] P+Dpred. 43.47 24.70 9.19 20.27 17.21 8.21 1.85 5.08 61.10 14.89 47.82 50.98 32.76 20.60
OccFormer [107] P 47.00 39.01 10.14 24.85 11.84 1.52 0.05 0.77 66.60 6.93 53.12 63.57 40.73 20.87
MonoScene [4] P 49.68 41.41 11.04 20.57 20.34 10.36 1.98 2.79 69.68 13.69 67.39 55.00 43.44 24.15